# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 5. A derived Milnor-Moore theorem
${\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
The symmetric monoidal forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ gives rise to a forgetful functor
${\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\subset\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\simeq\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
@@
Let ${\mathcal{C}}$ be a ${\mathbb{Q}}$-linear stable presentably symmetric monoidal $\infty$-category. The free ${\mathbb{E}}_{\infty}$-algebra functor
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{{\mathrm{Grp}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))}{\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$
@@
[…]
@@
[…]
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{Lie}}(1)}}\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nu}}_{\mathcal{C}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}),$
@@
${\mathcal{C}}\xrightarrow{{\mathrm{coComm}}^{\mathrm{nu}}\circ(-)}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nu}}_{\mathcal{C}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}_{\mathcal{C}}}({\mathcal{C}})\xrightarrow{(-)\oplus{\mathbb{1}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
[…]
${\mathcal{C}}\to{\mathcal{C}}:{\mathrm{X}}\mapsto\coprod_{{\mathrm{n}}\geq 0}({\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma_{\mathrm{n}}}$
@@
If ${\mathcal{C}}$ is not ${\mathbb{Q}}$-linear, we offer the following variant (Proposition 5.34) of Corollary 5.32. Let ${\mathcal{C}}$ be a preadditive presentably symmetric monoidal $\infty$-category. By Remark 2.31 the $\infty$-category ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})$ is presentable and thus admits finite products. So we can make the following definition:
Notation 5.33.
@@
${\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})).$
The forgetful functor ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ gives rise to a forgetful functor
@@
Let ${\mathcal{C}}$ be a stable presentably symmetric monoidal $\infty$-category that admits totalizations. The free ${\mathbb{E}}_{\infty}$-algebra functor
@@
By Proposition 5.4 the functor ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ preserves finite products and so gives rise to a functor
${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
@@
${\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
[…]
${\mathcal{C}}\to{\mathcal{C}}:{\mathrm{X}}\mapsto\coprod_{{\mathrm{n}}\geq 0}({\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma_{\mathrm{n}}}$
@@ 5.1. The classical Milnor-Moore theorem
(15) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})\rightleftarrows{\mathrm{Hopf}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}):{\mathrm{Prim}}$
@@
${\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}\to\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}),{\mathrm{X}}\mapsto\bigoplus_{{\mathrm{n}}\geq 0}({\mathrm{Lie}}_{\mathrm{n}}\otimes{\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma_{\mathrm{n}}}$
@@
(16) ${\mathrm{Prim}}_{\geq 0}:({\mathrm{coCoalg}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0})_{{\mathbb{1}}/}\simeq({\mathrm{coCoalg}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{{\mathbb{1}}}/)_{\geq 0}\to({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}$
sends augmented cocommutative coalgebra structures on ${\mathrm{K}}$-vector spaces to Lie algebras on ${\mathrm{K}}$-vector spaces. The functor (16) is right adjoint to the trivial cocommutative coalgebra functor $({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\to({\mathrm{coCoalg}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{{\mathbb{1}}/})_{\geq 0}$. The classical functor of primitive elements ${\mathrm{coCoalg}}(\mathrm{Vect}_{\mathrm{K}})_{{\mathbb{1}}/}\to\mathrm{Vect}_{{\mathrm{K}}}$ is right adjoint to the trivial cocommutative coalgebra functor $\mathrm{Vect}_{\mathrm{K}}\to{\mathrm{coCoalg}}(\mathrm{Vect}_{\mathrm{K}})_{{\mathbb{1}}/}$. By Künneth the embedding $\mathrm{Vect}_{\mathrm{K}}\hookrightarrow({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}$ is symmetric monoidal and admits a symmetric monoidal left adjoint $\pi_{0}:({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\to\mathrm{Vect}_{\mathrm{K}}$. Hence we obtain an induced adjunction
$\pi_{0}:{\mathrm{coCoalg}}(({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0})_{{\mathbb{1}}/}\rightleftarrows{\mathrm{coCoalg}}(\mathrm{Vect}_{\mathrm{K}})_{{\mathbb{1}}/}$
@@
[…]
@@
[…]
@@
$\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\simeq\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}(1)}({\mathcal{C}})\xrightarrow{\mathrm{TQ}_{{\mathrm{Ass}}_{\mathcal{C}}^{\mathrm{nu}}(1)}}{\mathrm{coAlg}}_{\mathrm{co}{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
[…]
@@
[…]
which is the functor $\nu\circ\theta_{!}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\xrightarrow{}{\mathcal{C}}$ since $\nu\circ\theta_{!}$ is symmetric monoidal and preserves sifted colimits and therefore preserves the Bar-construction. So it will be enough to see that the functor $\theta_{!}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\to\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}$ is symmetric monoidal and sends Lie structures on ${\mathrm{K}}$-vector spaces to associative algebra structures on ${\mathrm{K}}$-vector spaces. The canonical map $\sigma:{\mathrm{Lie}}_{\mathrm{K}}\to{\mathrm{Ass}}^{\mathrm{nu}}_{\mathrm{K}}$ of operads in chain complexes over ${\mathrm{K}}$ from the classical Lie operad over ${\mathrm{K}}$ to the non-unital associative operad over ${\mathrm{K}}$ is sent by the universal functor $\mathsf{Ch}_{\mathrm{K}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}$ to ${\mathrm{H}}({\mathrm{K}})\wedge\theta.$ The map $\sigma$ gives rise to a right adjoint forgetful functor
@@
We write ${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})\to\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}$ for the left adjoint of ${\mathrm{G}}$. The functor ${\mathcal{U}}^{\prime}$ is the classical enveloping algebra that sends ${\mathrm{L}}$ to the quotient of $\bigoplus_{{\mathrm{n}}\geq 0}{\mathrm{L}}^{\otimes{\mathrm{n}}}$ by the left and right ideal generated by the elements $[{\mathrm{A}},{\mathrm{B}}]-{\mathrm{A}}{\mathrm{B}}-{\mathrm{B}}{\mathrm{A}}$ for ${\mathrm{A}},{\mathrm{B}}\in{\mathrm{L}}.$ This description of ${\mathcal{U}}^{\prime}$ guarantees that ${\mathcal{U}}^{\prime}$ preserves quasi-isomorphisms since ${\mathrm{K}}$ is a field. The functor ${\mathcal{U}}^{\prime}$ is known to send the product to the tensor product. Thus ${\mathcal{U}}^{\prime}$ induces a symmetric monoidal functor
[…]

### v2

@@ 4. A derived Milnor-Moore theorem
${\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
The symmetric monoidal forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ gives rise to a forgetful functor
${\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\subset\mathrm{Mon}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\simeq\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
@@
[…]
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{{\mathrm{Grp}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))}{\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$
@@
${{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)))}$${{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))}$${{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))={\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).}$
@@
The composition ${\mathcal{C}}\xrightarrow{\phi}{\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\xrightarrow{\text{forget}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ is the functor
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{Lie}}(1)}}\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nun}}_{\mathcal{C}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}),$
@@
${\mathcal{C}}\xrightarrow{{\mathrm{coComm}}^{\mathrm{nun}}\circ(-)}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nun}}_{\mathcal{C}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nun}}_{\mathcal{C}}}({\mathcal{C}})\xrightarrow{(-)\oplus{\mathbb{1}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
Thus the composition $\phi^{\prime}:{\mathcal{C}}\xrightarrow{\phi}{\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\xrightarrow{\text{forget}}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ lifts the functor
${\mathcal{C}}\to{\mathcal{C}}:{\mathrm{X}}\mapsto\coprod_{{n}\geq 0}({\mathrm{X}}^{\otimes{n}})_{\Sigma_{n}}$
@@
If $\mathcal{C}$} is not $\mathbb{Q}$}-linear, we offer the following variant (Proposition 4.37) of Corollary 4.35. Let $\mathcal{C}$} be a preadditive presentably symmetric monoidal $\infty$-category. By Remark 2.31 the $\infty$-category $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nun}}}({\mathcal{C}})$ is presentable and thus admits finite products. So we can make the following definition:
Notation 4.36.
@@
${\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nun}}}({\mathcal{C}})).$
The forgetful functor $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}^{\mathrm{nun}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ gives rise to a forgetful functor
@@
[…]
@@
By Proposition 4.4 the functor ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ preserves finite products and so gives rise to a functor
${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
@@
${\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})={\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
By Remark 3.3 the composition $\phi^{\prime}:{\mathcal{C}}\xrightarrow{\phi}{\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\xrightarrow{\text{forget}}{\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\xrightarrow{\text{forget}}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ lifts the functor
${\mathcal{C}}\to{\mathcal{C}}:{\mathrm{X}}\mapsto\coprod_{{n}\geq 0}({\mathrm{X}}^{\otimes{n}})_{\Sigma_{n}}$
@@ 4.1. The classical Milnor-Moore theorem
(4.5) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})\rightleftarrows{\mathrm{Hopf}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}):{\mathrm{Prim}}$
@@
${\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}\to\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}),{\mathrm{X}}\mapsto\bigoplus_{{n}\geq 0}({\mathrm{Lie}}_{n}\otimes{\mathrm{X}}^{\otimes{n}})_{\Sigma_{n}}$
@@
(4.6) ${\mathrm{Prim}}_{\geq 0}:(\mathrm{co}\mathrm{Coalg}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0})_{{\mathbb{1}}/}\simeq(\mathrm{co}\mathrm{Coalg}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{{\mathbb{1}}}/)_{\geq 0}\to({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}$
sends augmented cocommutative coalgebra structures on $\mathrm{K}$}-vector spaces to Lie algebras on $\mathrm{K}$}-vector spaces. The functor (4.6) is right adjoint to the trivial cocommutative coalgebra functor $({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\to(\mathrm{co}\mathrm{Coalg}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{{\mathbb{1}}/})_{\geq 0}$. The classical functor of primitive elements $\mathrm{co}\mathrm{Coalg}({\mathrm{Vect}}_{\mathrm{K}})_{{\mathbb{1}}/}\to{\mathrm{Vect}}_{{\mathrm{K}}}$ is right adjoint to the trivial cocommutative coalgebra functor ${\mathrm{Vect}}_{\mathrm{K}}\to\mathrm{co}\mathrm{Coalg}({\mathrm{Vect}}_{\mathrm{K}})_{{\mathbb{1}}/}$. By Künneth the embedding ${\mathrm{Vect}}_{\mathrm{K}}\hookrightarrow({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}$ is symmetric monoidal and admits a symmetric monoidal left adjoint $\pi_{0}:({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\to{\mathrm{Vect}}_{\mathrm{K}}$. Hence we obtain an induced adjunction
$\pi_{0}:\mathrm{co}\mathrm{Coalg}(({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0})_{{\mathbb{1}}/}\rightleftarrows\mathrm{co}\mathrm{Coalg}({\mathrm{Vect}}_{\mathrm{K}})_{{\mathbb{1}}/}$
@@
${({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{0}}$$\scriptstyle{{\mathrm{triv}}}$${\mathrm{co}\mathrm{Coalg}(({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0})_{{\mathbb{1}}/}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{0}}$${{\mathrm{Vect}}_{\mathrm{K}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathrm{triv}}}$${\mathrm{co}\mathrm{Coalg}({\mathrm{Vect}}_{\mathrm{K}})_{{\mathbb{1}}/}.}$
@@
${\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\theta_{!}}$${\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\theta(1)_{!}}$$\scriptstyle{\mathrm{TQ}_{{\mathrm{Lie}}_{\mathcal{C}}(1)}}$${\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathrm{co}\mathrm{Coalg}({\mathcal{C}})_{{\mathbb{1}}/}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}(1)}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{TQ}_{{\mathrm{Ass}}_{\mathcal{C}}^{\mathrm{nu}}(1)}}$${\mathrm{co}\mathrm{Alg}_{\mathrm{co}{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}.}$
@@
$\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\simeq\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{Ass}}^{\mathrm{nu}}(1)}({\mathcal{C}})\xrightarrow{\mathrm{TQ}_{{\mathrm{Ass}}_{\mathcal{C}}^{\mathrm{nu}}(1)}}\mathrm{co}\mathrm{Alg}_{\mathrm{co}{\mathrm{Ass}}^{\mathrm{nu}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
$\rho:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\xrightarrow{\theta_{!}}\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\xrightarrow{\mathrm{Bar}}\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\text{forget}}{\mathcal{C}}$
@@
$\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}}))\xrightarrow{(\nu\circ\theta_{!})_{*}}\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}\xrightarrow{\mathrm{Bar}}\mathrm{co}\mathrm{Alg}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\text{forget}}{\mathcal{C}},$
which is the functor $\nu\circ\theta_{!}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\xrightarrow{}{\mathcal{C}}$ since $\nu\circ\theta_{!}$ is symmetric monoidal and preserves sifted colimits and therefore preserves the Bar-construction. So it will be enough to see that the functor $\theta_{!}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\to\mathrm{Alg}({\mathcal{C}})_{/{\mathbb{1}}}$ is symmetric monoidal and sends Lie structures on $\mathrm{K}$}-vector spaces to associative algebra structures on $\mathrm{K}$}-vector spaces. The canonical map $\sigma:{\mathrm{Lie}}_{\mathrm{K}}\to{\mathrm{Ass}}^{\mathrm{nu}}_{\mathrm{K}}$ of operads in chain complexes over $\mathrm{K}$} from the classical Lie operad over $\mathrm{K}$} to the non-unital associative operad over $\mathrm{K}$} is sent by the universal functor $\mathsf{Ch}_{\mathrm{K}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}$ to ${\mathrm{H}}({\mathrm{K}})\smash{\theta}.$ The map $\sigma$ gives rise to a right adjoint forgetful functor
@@
We write ${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})\to\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}$ for the left adjoint of $\mathrm{G}$}. The functor $\mathcal{U}$$^{\prime}$ is the classical enveloping algebra that sends $\mathrm{L}$} to the quotient of $\bigoplus_{{n}\geq 0}{\mathrm{L}}^{\otimes{n}}$ by the left and right ideal generated by the elements ${\mathrm{A}},{\mathrm{B}}]-{\mathrm{A}}{\mathrm{B}}-{\mathrm{B}}{\mathrm{A}}$ for ${\mathrm{A}},{\mathrm{B}}\in{\mathrm{L}}.$ This description of $\mathcal{U}$$^{\prime}$ guarantees that $\mathcal{U}$$^{\prime}$ preserves quasi-isomorphisms since $\mathrm{K}$} is a field. The functor $\mathcal{U}$$^{\prime}$ is known to send the product to the tensor product. Thus $\mathcal{U}$$^{\prime}$ induces a symmetric monoidal functor
(4.7) $\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})[\text{quasi-isos}]\to\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}[\text{quasi-isos}]$